The projection of the line segment joining (1,2,-1), (4,2,1) on the line x/2=-y=z is 7/2
True or false with correct explanation
1
Expert's answer
2020-11-17T16:41:00-0500

By definition, the projection of the line segment joining "(1,2,-1), (4,2,1)" on the line "\frac{x}{2}=\frac{y}{-1}=\frac{z}{1}" is a absolute value of the projection of the vector "\overline{a}=(4,2,1)-(1,2,-1)=(3,0,2)" on the vector "\overline{b}=(2,-1,1)" which is parallel to the above line. Therefore, "pr_{\overline{b}}\overline a=\frac{|\overline{a}\cdot\overline{b}|}{|\overline{b}|} = \frac{|3\cdot 2+0\cdot(-1)+2\cdot 1|}{\sqrt{4+1+1}}=\frac{8}{\sqrt{6}}."

We conclude that it is not true that the projection of the line segment joining "(1,2,-1), (4,2,1)" on the line "\frac{x}{2}=\frac{y}{-1}=\frac{z}{1}" is "\frac{7}{2}".